# Self-supervised Learning of LiDAR 3D Point Clouds via 2D-3D Neural Calibration

### v1

@@ Abstract
This paper introduces a novel self-supervised learning framework for enhancing 3D perception in autonomous driving scenes. Specifically, our approach, named NCLR, focuses on 2D-3D neural calibration, a novel pretext task that estimates the rigid transformation aligning camera and LiDAR coordinate systems. First, we propose the learnable transformation alignment to bridge the domain gap between image and point cloud data, converting features into a unified representation space for effective comparison and matching. Second, we identify the overlapping area between the image and point cloud with the fused features. Third, we establish dense 2D-3D correspondences to estimate the rigid transformation. The framework not only learns fine-grained matching from points to pixels but also achieves alignment of the image and point cloud at a holistic level, understanding their relative pose. We demonstrate NCLR’s efficacy by applying the pre-trained backbone to downstream tasks, such as LiDAR-based 3D semantic segmentation, object detection, and panoptic segmentation. Comprehensive experiments on various datasets illustrate the superiority of NCLR over existing self-supervised methods. The results confirm that joint learning from different modalities significantly enhances the network’s understanding abilities and effectiveness of learned representation. Code will be available at https://github.com/Eaphan/NCLR.
@@
Self-supervised Learning, 3D Perception, Cross-modal, Autonomous Driving.
1 Introduction
The LiDAR technology serves as a vital enhancement to 2D cameras by precisely capturing the surroundings of a vehicle, offering robust performance in challenging conditions, including low light, intense sunlight, or glare from approaching headlights. This 3D perception, derived from LiDAR point clouds, is essential for the effective navigation of autonomous vehicles in three-dimensional spaces. While current leading methods are based on extensive labeled datasets, the process of labeling 3D data is exceedingly expensive and time-intensive, given the limited availability of annotation resources. Consequently, there is a growing imperative to utilize unlabeled data. This approach aims to minimize the need for extensive annotation while enhancing the effectiveness of networks trained on a limited amount of labeled data.
@@ 1 Introduction
Figure 1: Illustration of the proposed 2D-3D neural calibration developed as the pretext task for self-supervised learning of 3D LiDAR point clouds. Our method not only learns the local matching from points to pixels via contrastive losses, but also estimates the (unknown) holistic rigid transformation between camera and LiDAR systems. Besides, we propose learnable transformation alignment to fill the domain gap between the image and the point cloud.
@@
In this paper, we introduce a novel pretext task, the 2D-3D neural calibration, for the self-supervised pre-training of networks for 3D perception in autonomous driving scenes, as shown in Fig. 1. During pre-training, our pretext task not only learns fine-grained matching from individual points to pixels but also achieves a comprehensive alignment between the image and point cloud data, i.e., understanding their relative pose. Specifically, to impose supervision on the rigid transformation estimation, we propose an end-to-end differentiable framework distinguished by its integration of a soft-matching strategy and a differentiable PnP solver. We posit that this joint learning from different modalities for both local and global-level alignment will enhance the network’s capacity for sophisticated understanding and enable it to develop effective representations. In addition, to fill the substantial domain gap between point clouds and images, we propose a learnable transformation alignment during pre-training, replacing the direct cosine distance alignment. This method converts features into a unified representation space, allowing for more accurate feature comparison and matching beyond the capabilities of cosine distance.
@@
The rest of this paper is organized as follows. Section 2 provides an overview of related literature pertinent to our research. In Section 3, we present the overall architecture of NCLR and elaborate its principal elements. Section 4 presents an empirical evaluation of our proposed approach across three distinct downstream tasks, along with ablation studies to assess the impact of key components. The paper concludes with Section 5, summarizing our findings.
2 Related Work
2.1 Self-supervised Learning for Images
@@
Self-supervised learning involves training models on unlabeled data by generating pseudo-labels or tasks based on the data itself, without human annotations. The model learns representations by solving pretext problems derived from the inherent structure of the unlabeled inputs. The field of self-supervised learning on images has seen significant evolution over the years. Initial approaches were centered around pretext tasks, which are auxiliary tasks designed to derive meaningful data representations. These tasks involved teaching models to restore color in black-and-white images [23] and to ascertain the spatial relation between image segments [24].
@@
The advent of deep learning brought about a paradigm shift in the field. One such approach is contrastive learning [25, 26]. Contrastive learning involves training networks to distinguish between different augmentations of the same image versus those from distinct images. It minimizes the representation distance for augmentations of the same image (positive pairs) while maximizing it for different images (negative pairs). Generative models [2, 27] have also emerged as a significant trend in self-supervised learning. These models, often employing Variational Autoencoders (VAEs) or Generative Adversarial Networks (GANs), learned to generate images that closely resemble the training data. By learning to generate images, these models captured the underlying data distribution, which can be leveraged for various downstream tasks. In addition to these, there have been several exciting advancements in the field. Masked image modeling, for instance, involves masking parts of an image and training a model to predict the masked parts [28]. Multi-modal models aim to learn representations across different modalities (e.g., text, image, audio) using self-supervised learning. The Vision Transformer (ViT) adapted the transformer architecture, initially developed for natural language processing, to computer vision tasks [29]. Additionally, there has been explored into incorporating physics-based priors in self-supervised learning.
2.2 Self-supervised Learning for Point Clouds
@@
The advances in self-supervision on point clouds have closely followed the improvements made in images. Early self-supervised methods used pretext tasks such as predicting transformations applied to the point cloud or reconstructing parts of the point cloud [30, 31]. These methods were applied on dense scans of single objects [32, 33].
@@
Discriminative-based Methods. This type of method explored different levels of representation (point-level [3, 4], segment-level [5], region-level [6, 7]) and introduced contrastive losses to capture the geometric and structural nuances of 3D data [8, 9]. Nunes et al. introduced a novel contrastive learning approach focused on understanding scene structure. This method involves extracting class-agnostic segments from point clouds and employing contrastive loss on these segments to differentiate between structurally similar and dissimilar elements [5].
@@
Temporal-consistency-based Methods. STRL [11] utilized pairs of temporally-linked frames from 3D point cloud sequences, applying spatial data augmentation to learn invariant representations. STSSL [12] incorporated spatial-temporal positive pairs, introducing a point-to-cluster technique for spatial object distinction and a cluster-to-cluster method using unsupervised tracking for temporal correlations. TARL [10] used vehicle motion to align objects over time in different scans, focusing on enhancing segment-level feature similarities for consistent temporal representations. These works demonstrated how leveraging temporal information can lead to robust and invariant representations, which are beneficial for various downstream tasks.
@@
Reconstruction-based Methods. Reconstruction-based methods have also been successful for self-supervision on point clouds. Some methods reconstructed point coordinates using the Chamfer distance [13, 14]. Point-BERT randomly masked some patches and fed them into Transformers to recover the original tokens at the masked locations [34]. Recently, ALSO [15] proposed to use unsupervised surface reconstruction as a pretext task to train 3D backbones on automotive LiDAR point clouds. Using the knowledge of occupancy before and after an observed 3D point along a LiDAR ray, it learned to construct an implicit occupancy function and good point features.
@@
Multi-modal Self-supervised Learning Another line of work leveraged synchronized and calibrated cameras and LiDAR to pretrain a 3D backbone. The underlying idea is to find pairs of corresponding points and pixels and ensure that the associated point and pixel representations are as close as possible [16, 17, 18].
2.3 LiDAR-based 3D Perception
@@
3D Semantic Segmentation. One of the foundational works in this area is PointNet [36], which marked a significant departure from previous methods by directly processing point clouds for tasks like segmentation. MinkUNet adapted the U-Net architecture for sparse 3D point clouds, using Minkowski Engine for efficient segmentation [37]. SPVCNN combined sparse and continuous convolutions for scalable and efficient 3D point cloud segmentation [38].
3D Object Detection. Advancements in 3D object detection have led to its categorization into voxel-based, point-based, and hybrid methods [39, 40, 41, 42]. Voxel-based techniques convert point clouds into uniform grid structures for feature extraction using sparse convolutions [43, 44]. Point-based methods directly handle raw 3D point clouds to make predictions. Hybrid models, such as PV-RCNN [45] and PDV [46], merge the advantages of both voxel and point-based approaches, optimizing both accuracy and efficiency in detection.
3D Panoptic Segmentation. The task requires a holistic understanding of both objects and scenes. Hong et al. introduced the Dynamic Shifting Network (DS-Net) as a potent framework for panoptic segmentation in the field of point clouds [47]. This approach utilized cylinder convolution tailored for LiDAR data. The features derived were concurrently used by both the semantic and instance branches, functioning through a bottom-up clustering approach.
3 Proposed Method
3.1 Overview
We propose a novel approach to generating self-supervised features for 3D point clouds, utilizing the 2D-3D neural calibration as a pretext task. The 2D-3D neural calibration aims to identify the unknown rigid transformation, comprising a rotation $\mathbf{R}\in SO(3)$ and a translation $\mathbf{t}\in\mathbb{R}^{3}$, through a neural network. This transformation aligns the coordinate systems of the camera and the point cloud. We denote the input image as $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$ and point cloud as $\mathbf{P}=\{\mathbf{p}_{i}\in\mathbb{R}^{3}|i=1,2,...,N\}$, where $H$ and $W$ are the height and width of the image, and $N$ is the number of points. To achieve effective representations, our pre-training method not only learns fine-grained matching from points to pixels but also achieves alignment of the image and point cloud at a holistic level, i.e., understanding their relative pose. Then the point cloud backbone could be used for downstream 3D perception tasks.
@@ 3.3 Feature Discrimination
We illustrate the learned features using cosine distance and learnable transformation alignment, respectively, in Fig. 4. We can observe that it is non-trivial to drive the matching 2D and 3D features to be similar with respect to cosine distance due to the inherent domain gap between point clouds and images. In contrast, when we use the learnable transformation alignment for computing similarities during training, the 2D and 3D features are closer and evenly distributed in the transformed unified representation space. We further analyze the effect of learnable transformation alignment in Section 4.5.
3.4 Building Dense Correspondence
@@
3.5 Estimation of Rigid Transformation
The final module in our framework estimates the rigid transformation based on the soft matching results computed in Section 3.4. The neural calibration problem is defined as determining the rigid transformation $\{\mathbf{R},\mathbf{t}\}$ that best maps points in overlapping areas $\mathbf{P}_{o}$ onto the estimated 2D coordinates in the image plane, i.e.,
$\widehat{\mathbf{R}},\widehat{\mathbf{t}}=\mathop{\arg\min}_{\mathbf{R},\mathbf{t}}\sum_{\mathbf{p}_{i}\in\mathbf{P}_{o}}\left[\mathbf{K}(\mathbf{R}\mathbf{p}_{i}+\mathbf{t})-\widehat{\mathbf{C}}_{i}^{\mathbf{P}}\right].$ (6)
In this paper, we solve Eq. (6) using a differentiable EPnP algorithm [51].
The rigid transformation estimation is optimized by the loss function defined as:
@@ 3.5 Estimation of Rigid Transformation
3.6 Training Pipeline
Initially, each pair of an image and a point cloud is associated with a relative pose, denoted as $\{\mathbf{R}_{raw},\mathbf{t}_{raw}\}$. Since the camera and LiDAR sensors are fixed in a relatively close position of the cars, the original transformation from the camera coordinate to the point cloud coordinate is similar. To avoid the network overfitting to such scenarios, we simulate a range of sensor placements and orientations by applying random rotation $\mathbf{R}_{r}$ and translation $\mathbf{t}_{r}$ to the original point cloud during the pre-training process. The ground-truth pose is recalculated to accommodate these adjustments:
@@ 3.6 Training Pipeline
4 Experiments
In this section, we begin with an overview of the datasets and evaluation metrics in Sec.4.1. Following this, Sec.4.2 to Sec.4.4 detail the experimental setup and the fine-tuning results for three downstream tasks. Subsequent sections, Sec.4.5 and Sec.4.6, delve into comprehensive ablation studies and in-depth analyses to assess critical aspects of our framework. Finally, we explore the reasons why the pre-text task is effective in Sec 4.8.
4.1 Dataset
@@
SemanticKITTI (SK) dataset contains RGB image and point cloud data pairs from KITTI scenes for the task of urban scene semantic segmentation [19]. The data is collected from sensor systems mounted on a car, comprising over 200,000 images and corresponding point clouds from 21 different scenes/sequences. The images and point clouds are synchronized to ensure a fixed relative transformation. The original images have a resolution of 1241x376 pixels. Each point cloud contains approximately 40,000 3D points. Following common practice, the dataset is split into a training set using the 10 sequences and a validation set using the eighth sequence.
TABLE I: Results of semantic segmentation models fine-tuned on three distinct datasets. This comparison considers the quantity of annotated data, the dataset employed for pre-training, and variations in model architecture. Our analysis contrasts NCLR with a baseline that lacks pre-training and other contemporary self-supervised methods, using the mean Intersection over Union (mIoU) percentage metric.
@@
4.2 Semantic Segmentation

### v2

@@ Abstract
This paper introduces a novel self-supervised learning framework for enhancing 3D perception in autonomous driving scenes. Specifically, our approach, namely NCLR, focuses on 2D-3D neural calibration, a novel pretext task that estimates the rigid pose aligning camera and LiDAR coordinate systems. First, we propose the learnable transformation alignment to bridge the domain gap between image and point cloud data, converting features into a unified representation space for effective comparison and matching. Second, we identify the overlapping area between the image and point cloud with the fused features. Third, we establish dense 2D-3D correspondences to estimate the rigid pose. The framework not only learns fine-grained matching from points to pixels but also achieves alignment of the image and point cloud at a holistic level, understanding their relative pose. We demonstrate the efficacy of NCLR by applying the pre-trained backbone to downstream tasks, such as LiDAR-based 3D semantic segmentation, object detection, and panoptic segmentation. Comprehensive experiments on various datasets illustrate the superiority of NCLR over existing self-supervised methods. The results confirm that joint learning from different modalities significantly enhances the network’s understanding abilities and effectiveness of learned representation. The code is publicly available at https://github.com/Eaphan/NCLR.
@@
Self-supervised Learning, 3D Perception, Cross-modal, Autonomous Driving, Registration.
I Introduction
The LiDAR technology serves as a vital enhancement to 2D cameras by precisely capturing the surroundings of a vehicle, offering robust performance in challenging conditions, including low light, intense sunlight, or glare from approaching headlights. This 3D perception, derived from LiDAR point clouds, is essential for the effective navigation of autonomous vehicles in three-dimensional spaces. While current leading methods are based on extensive labeled datasets, labeling 3D data is exceedingly expensive and time-intensive, given the limited availability of annotation resources. Consequently, there is a growing imperative to utilize unlabeled data. This approach aims to minimize the need for extensive annotation while enhancing the effectiveness of networks trained on a limited amount of labeled data.
@@ I Introduction
Figure 1: Illustration of the proposed 2D-3D neural calibration developed as the pretext task for self-supervised learning of 3D LiDAR point clouds. Our method not only learns the local matching from points to pixels via contrastive losses but also estimates the (unknown) holistic rigid pose aligning the camera and LiDAR systems. Besides, we propose learnable transformation alignment to fill the domain gap between the image and the point cloud.
@@
In this paper, we introduce a novel pretext task, the 2D-3D neural calibration, for the self-supervised pre-training of networks for 3D perception in autonomous driving scenes, as shown in Fig. 1. During pre-training, our pretext task not only learns fine-grained matching from individual points to pixels but also achieves a comprehensive alignment between the image and point cloud data, i.e., understanding their relative pose. Specifically, to impose supervision on the rigid pose estimation, we propose an end-to-end differentiable framework distinguished by its integration of a soft-matching strategy and a differentiable PnP solver. We posit that this joint learning from different modalities for both local and global-level alignment will enhance the network’s capacity for sophisticated understanding and enable it to develop effective representations. In addition, to fill the substantial domain gap between point clouds and images, we propose a learnable transformation alignment during pre-training, replacing the direct cosine distance alignment. This method converts features into a unified representation space, allowing for more accurate feature comparison and matching beyond the capabilities of cosine distance.
@@
The rest of this paper is organized as follows. Section II provides an overview of related literature pertinent to our research. In Section III, we present the overall architecture of NCLR and elaborate its principal elements. Section IV presents an empirical evaluation of our proposed approach across three distinct downstream tasks, along with ablation studies to assess the impact of key components. The paper concludes with Section V, summarizing our findings.
II Related Work
II-A Self-supervised Learning for Images
@@
Self-supervised learning involves training models on unlabeled data by generating pseudo-labels or tasks based on the data itself, without human annotations. The model learns representations by solving pretext problems derived from the inherent structure of the unlabeled inputs. The field of self-supervised learning on images has seen significant evolution over the years. Initial approaches are centered around pretext tasks, which are auxiliary tasks designed to derive meaningful data representations. These tasks involve teaching models to restore color in black-and-white images [23] and to ascertain the spatial relation between image segments [24].
@@
The advent of deep learning has brought about a paradigm shift in the field. One such approach is contrastive learning [25, 26]. Contrastive learning involves training networks to distinguish between different augmentations of the same image versus those from distinct images. It minimizes the representation distance for augmentations of the same image (positive pairs) while maximizing it for different images (negative pairs). Generative models [2, 27] have also emerged as a significant trend in self-supervised learning. These models, often employing Variational Autoencoders (VAEs) or Generative Adversarial Networks (GANs), learn to generate images that closely resemble the training data. By learning to generate images, these models capture the underlying data distribution, which can be leveraged for various downstream tasks. In addition to these, there have been several exciting advancements in the field. Masked image modeling, for instance, involves masking parts of an image and training a model to predict the masked parts [28]. Multi-modal models aim to learn representations across different modalities (e.g., text, image, audio) using self-supervised learning. The Vision Transformer (ViT) adapts the transformer architecture, initially developed for natural language processing, to computer vision tasks [29]. Additionally, there has been explored into incorporating physics-based priors in self-supervised learning.
II-B Self-supervised Learning for Point Clouds
@@
The advances in self-supervision on point clouds have closely followed the improvements made in images. Early self-supervised methods use pretext tasks such as predicting transformations applied to the point cloud or reconstructing parts of the point cloud [30, 31]. These methods are applied on dense scans of single objects [32, 33].
@@
Discriminative-based Methods. This type of method explores different levels of representation (point-level [3, 4], segment-level [5], region-level [6, 7]) and introduces contrastive losses to capture the geometric and structural nuances of 3D data [8, 9]. Nunes et al. introduced a novel contrastive learning approach focused on understanding scene structure. This method involves extracting class-agnostic segments from point clouds and employing contrastive loss on these segments to differentiate between structurally similar and dissimilar elements [5].
@@
Temporal-consistency-based Methods. STRL [11] utilizes pairs of temporally-linked frames from 3D point cloud sequences, applying spatial data augmentation to learn invariant representations. STSSL [12] incorporates spatial-temporal positive pairs, introducing a point-to-cluster technique for spatial object distinction and a cluster-to-cluster method using unsupervised tracking for temporal correlations. TARL [10] uses vehicle motion to align objects over time in different scans, focusing on enhancing segment-level feature similarities for consistent temporal representations. These works demonstrate how leveraging temporal information can lead to robust and invariant representations, which are beneficial for various downstream tasks.
@@
Reconstruction-based Methods. Reconstruction-based methods have also been successful for self-supervision on point clouds. Some methods reconstruct point coordinates using the Chamfer distance [13, 14]. Point-BERT randomly masks some patches and feeds them into Transformers to recover the original tokens at the masked locations [34]. Recently, ALSO [15] proposes to use unsupervised surface reconstruction as a pretext task to train 3D backbones on automotive LiDAR point clouds. Using the knowledge of occupancy before and after an observed 3D point along a LiDAR ray, it learns to construct an implicit occupancy function and good point features. Besides, rendering-based techniques [35, 36, 37] reconstruct masked scenes from point clouds or multi-view images, which have achieved promising generalization ability.
@@
Multi-modal Self-supervised Learning Another line of work leverages synchronized and calibrated cameras and LiDAR to pretrain a 3D backbone. These contrastive methods find pairs of corresponding points and pixels and ensure that the associated point and pixel representations are as close as possible [16, 17, 18]. The primary motivation of these methods is to transfer the knowledge already embedded in a pre-trained 2D backbone to the 3D backbone. In contrast, our approach trains the network to achieve holistic alignment between the image and point cloud, enabling it to understand the two modalities thoroughly and subsequently learn effective 3D representations. Besides, our method does not rely on a pre-trained 2D backbone. By eliminating this dependency, our approach introduces greater flexibility and removes the constraint of inheriting specific representations from a 2D network.
II-C LiDAR-based 3D Perception
@@
3D Semantic Segmentation. One of the foundational works in this area is PointNet [39], which marks a significant departure from previous methods by directly processing point clouds for tasks like segmentation. MinkUNet adapts the U-Net architecture for sparse 3D point clouds, using Minkowski Engine for efficient segmentation [40]. SPVCNN combines sparse and continuous convolutions for scalable and efficient 3D point cloud segmentation [41].
3D Object Detection. Advancements in 3D object detection have led to its categorization into voxel-based, point-based, and hybrid methods [42]. Voxel-based techniques convert point clouds into uniform grid structures for feature extraction using sparse convolutions [43, 44]. Point-based methods directly handle raw 3D point clouds to make predictions. Hybrid models, such as PV-RCNN [45] and PDV [46], merge the advantages of both voxel and point-based approaches, optimizing both accuracy and efficiency in detection.
3D Panoptic Segmentation. The task requires a holistic understanding of both objects and scenes. Hong et al. introduced the Dynamic Shifting Network (DS-Net) as a potent framework for panoptic segmentation in the field of point clouds [47]. This approach utilizes cylinder convolution tailored for LiDAR data. The features derived are concurrently used by both the semantic and instance branches, functioning through a bottom-up clustering approach.
III Proposed Method
III-A Overview
We propose a novel approach to generating self-supervised features for 3D point clouds, utilizing the 2D-3D neural calibration as a pretext task. The 2D-3D neural calibration aims to identify the unknown rigid pose, comprising a rotation $\mathbf{R}\in SO(3)$ and a translation $\mathbf{t}\in\mathbb{R}^{3}$, through a neural network. This pose aligns the coordinate systems of the camera and the point cloud. We denote the input image as $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$ and point cloud as $\mathbf{P}=\{\mathbf{p}_{i}\in\mathbb{R}^{3}|i=1,2,...,N\}$, where $H$ and $W$ are the height and width of the image, and $N$ is the number of points. To achieve effective representations, our pre-training method not only learns fine-grained matching from points to pixels but also achieves alignment of the image and point cloud at a holistic level, i.e., understanding their relative pose. Then the point cloud backbone could be used for downstream 3D perception tasks.
@@ III-C Feature Discrimination
We illustrate the learned features using cosine distance and learnable transformation alignment, respectively, in Fig. 4. We can observe that it is non-trivial to drive the matching 2D and 3D features to be similar with respect to cosine distance due to the inherent domain gap between point clouds and images. In contrast, when we use the learnable transformation alignment for computing similarities during training, the 2D and 3D features are closer and evenly distributed in the transformed unified representation space. We further analyze the effect of learnable transformation alignment in Section IV-E.
III-D Building Dense Correspondence
@@
III-E Estimation of Rigid Pose
The final module in our framework estimates the rigid pose based on the soft matching results computed in Section III-D. The neural calibration problem is defined as determining the rigid pose $\{\mathbf{R},\mathbf{t}\}$ that best maps points in overlapping areas $\mathbf{P}_{o}$ onto the estimated 2D coordinates in the image plane, i.e.,
$\widehat{\mathbf{R}},\widehat{\mathbf{t}}=\mathop{\arg\min}_{\mathbf{R},\mathbf{t}}\sum_{\mathbf{p}_{i}\in\mathbf{P}_{o}}\left[\mathbf{K}(\mathbf{R}\mathbf{p}_{i}+\mathbf{t})-\widehat{\mathbf{C}}_{i}^{\mathbf{P}}\right],$ (6)
where $\mathbf{K}$ denotes the camera intrinsic matrix.
In this paper, we solve Eq. (6) using a differentiable EPnP algorithm [51]. The rigid pose estimation is optimized by the loss function defined as:
@@ III-E Estimation of Rigid Pose
III-F Training Pipeline
Initially, each pair of an image and a point cloud is associated with a relative pose, denoted as $\{\mathbf{R}_{raw},\mathbf{t}_{raw}\}$. Since the camera and LiDAR sensors are fixed in a relatively close position of the cars, the original pose aligning the camera coordinate to the point cloud coordinate is similar. To avoid the network overfitting to such scenarios, we simulate a range of sensor placements and orientations by applying random rotation $\mathbf{R}_{r}$ and translation $\mathbf{t}_{r}$ to the original point cloud during the pre-training process. The ground-truth pose is recalculated to accommodate these adjustments:
@@ III-F Training Pipeline
IV Experiments
In this section, we begin with an overview of the datasets and evaluation metrics in Sec.IV-A. Following this, Sec.IV-B to Sec.IV-D detail the experimental setup and the fine-tuning results for three downstream tasks. Subsequent sections, Sec.IV-E and Sec.IV-F, delve into comprehensive ablation studies and in-depth analyses to assess critical aspects of our framework. Finally, we explore the reasons why the pre-text task is effective in Sec IV-H.
IV-A Dataset
@@
SemanticKITTI (SK) dataset contains RGB image and point cloud data pairs from KITTI scenes for the task of urban scene semantic segmentation [19]. The data is collected from sensor systems mounted on a car, comprising over 200,000 images and corresponding point clouds from 21 different scenes/sequences. The images and point clouds are synchronized to ensure a fixed relative pose. The original images have a resolution of 1241x376 pixels. Each point cloud contains approximately 40,000 3D points. Following common practice, the dataset is split into a training set using the 10 sequences and a validation set using the eighth sequence.
TABLE I: Results of semantic segmentation models fine-tuned on three distinct datasets. This comparison considers the quantity of annotated data, the dataset employed for pre-training, and variations in model architecture. Our analysis contrasts NCLR with a baseline that lacks pre-training and other contemporary self-supervised methods, using the mean Intersection over Union (mIoU) percentage metric. The best results are highlighted in bold.
@@
TABLE II: Comparison of various pre-training techniques for semantic segmentation tasks using either finetuning or linear probing. We report the mean Intersection over Union (mIoU) on the nuScenes validation set across different annotation levels.
Initialization Reference nuScenes SemanticKITTI
Lin. Prob Finetune Finetune
100% 1% 1%
Random - 8.10 30.30 39.50
PPKT [16] Arxiv’21 35.90 37.80 44.00
SLidR [17] CVPR’22 38.80 38.30 44.60
ST-SLidR [18] CVPR’23 40.48 40.75 44.72
Ours - 42.81 42.29 45.84
IV-B Semantic Segmentation